# Descent Methods for Vector Optimization Problems: A Majorization-Minimization Perspective

### v2

@@ 4.2 Improved $K$-steepest descent method for VOPs without line search
As detailed in [8], the linear convergence result in Lemma 4.1 can be improved, this is mainly due to the large gap between $F(\cdot)-F(x^{k})$ and $G_{k,Le}(\cdot)$ from a majorization-minimization perspective. Notice that $\bm{\ell}\preceq_{K}L_{\max}e$, we denote the following tighter surrogate:
@@
    $G_{k,\bm{\ell}}(\cdot)$ is a tight majorizing surrogate of $F(\cdot)-F(x^{k})$, i.e., $F(\cdot)-F(x^{k})\preceq_{K}G_{k,\bm{\ell}}(\cdot)$, and the relation does not hold for any $G_{k,\hat{\bm{\ell}}}(\cdot)$ such that $\bm{\ell}\not\preceq_{K}\hat{\bm{\ell}}$.
@@
By using the tighter surrogate, we devise the following improved $K$-steepest descent method without line search.
@@
Algorithm 3 #1_improved_$K$-steepest_descent_method_without_line_search_for_VOPs
@@
If $K=\mathbb{R}^{m}_{+}$, and $e=(1,...,1)^{T}$, the convergence rate in Lemma 4.2 reduces to that of [8, Corollary 4.3] with $g(\cdot)=0$. Furthermore, $1/\kappa_{F,\preceq_{K}}\geq\mu_{\min}/L$, and the improved linear convergence does not depend on the choice of $e$.
4.3 Equivalence between Algorithms 2 and 3
As we described in Remark 4.2, the linear convergence rate of Algorithms 2 depends on $e$. In response to this issue, we attempt to give a better choice of $e$ in Algorithms 2 from a complexity perspective. Recall that the linear convergence rate of algorithm 3 does not depend on the choice of $e$. We denote
$C_{\bm{\ell}}:=\{c^{*}\in K^{*}:\left\langle{c^{*},\bm{\ell}}\right\rangle=1\}.$
In the following, we propose another improved $K$-steepest descent method for VOPs.
@@ 4.3 Equivalence between Algorithms 2 and 3
Algorithm 4 #2_improved_$K$-steepest_descent_method_without_line_search_for_VOPs
Notice that the subproblem in Algorithm 4 can be rewritten equivalently as:
@@
Similar to Lemma 4.2, Algorithm 4 enjoys the following improved linear convergence.
@@
If $K=\mathbb{R}^{m}_{+}$, and $\Delta_{m}^{\bm{\ell}}:=\{c^{*}\in\mathbb{R}^{m}_{++}:\left\langle{c^{*},\bm{\ell}}\right\rangle=1\}$, the Algorithm 4 reduces to [8, Algorithm 5] with $g(\cdot)=0$. Interestingly, the relations between Algorithms 2, 3 and 4 only depends on the choice of $e$. If $e=\bm{\ell}$, Algorithms 2, 3 and 4 are equivalent. As a result, for the open problem of the better choice of $e$ in Remark 4.2, we give a partially positive answer with $e=\bm{\ell}$.
@@
Although Algorithms 3 and 4 both exhibit similar improved linear convergence, the computational cost of solving subproblems is generally lower in Algorithm 4. Details on solving these subproblems will be provided in Section 6.
5 Descent methods for VOPs with line search
@@ 5.1 $K$-steepest descent method with line search
In this section, we revisit $K$-steepest descent method with line search for VOPs.
@@
4             return $C$-stationary point $x^{k}$
@@ 7             $x^{k+1}:=x^{k}+t_{k}d^{k}$
Algorithm 5 K-steepest_descent_method_for_VOPs
@@
We can show that $G_{k,e/t_{\min}}(\cdot)$ is a non-majorizing surrogate function of $F(\cdot)-F(x^{k})$ near $x^{k}$.
@@
If $K=\mathbb{R}^{m}_{+}$, and $e=(1,...,1)^{T}$, the convergence rate in Lemma 5.1(ii) reduces to that of [19, Theorem 4.2].
5.2 Generic first-order method for VOPs
Select $e_{k}\in{\rm int}(K)$, we devise the following generic first-order method:
@@ 2      Select $e_{k}\in{\rm int}(K)$
5             return $C$-stationary point $x^{k}$
@@ 8             $x^{k+1}:=x^{k}+t_{k}d^{k}$
Algorithm 6 Generic_first_order_method_for_VOPs
It is worth noting that we don’t specify how to select $e_{k}$ in Algorithm 6. This naturally raises the question: what role does $e_{k}$ play in determining the convergence rate?
Firstly, we derive the lower bound of stepsize in each iteration.
@@
    $\{x^{k}\}$ converges to an efficient solution $x^{*}$ of (VOP);
@@
As described in Lemma 5.2(iv), by setting $e_{k}=\bm{\mu}$ or $e_{k}=\bm{\ell}$, we can derive the optimal linear convergence rate for Algorithm 6. Moreover, the convergence rate reduces to that of Lemma 4.3(ii) with constant $\gamma$. Intuitively, to explore the local curvature information of $F(\cdot)$, we can devise a tighter local surrogate $G_{k,e_{k}/t_{k}}(\cdot)$ with $\bm{\mu}\preceq_{K}e_{k}\preceq_{K}\bm{\ell}$. In this case, the linear convergence rate of Algorithm 6 can be further improved by using a tighter local surrogate $G_{k,e_{k}/t_{k}}(\cdot)$.
@@
6 Descent methods for VOPs with polyhedral cones
@@ 6.1 Steepest descent method for VOPs with polyhedral cones
By using the transform matrix $A\in\mathcal{A}$, the steepest descent direction subproblem for VOPs with polyhedral cones is formulated as follows:
@@
In the next subsection, we claim that the preceding subproblem can be efficiently solved via its duality. The complete $K$-steepest descent method for VOPs with polyhedral cones is described as follows:
@@ 3       Update $d^{k}$ as the minimizer of (16)
5             return $C$-stationary point $x^{k}$
@@ 7            Compute the stepsize $t_{k}\in(0,1]$ in the following way:
$t_{k}:=\max\left\{\gamma^{j}:j\in\mathbb{N},A(F(x^{k}+\gamma^{j}d^{k})-F(x^{k}))\preceq\gamma^{j}\left(AJF(x^{k})d^{k}+\frac{1}{2}\left\lVert{d^{k}}\right\rVert^{2}e\right)\right\},$
where $e:=(1,...,1)^{T}$
8             $x^{k+1}:=x^{k}+t_{k}d^{k}$
@@
Algorithm 7 K-steepest_descent_method_for_VOPs_with_polyhedral_cones
@@
    $t_{k}\geq\min_{i\in[l]}\{\gamma/\left\langle{A_{i},\bm{\ell}}\right\rangle\}$;
@@
    $\left\lVert{x^{k+1}-x^{*}}\right\rVert\leq\sqrt{1-\min_{i\in[l]}t_{k}\left\langle{A_{i},\bm{\mu}}\right\rangle}\left\lVert{x^{k}-x^{*}}\right\rVert,~{}\forall k\geq 0$.
@@ 6.2 Barzilai-Borwein descent method for VOPs with polyhedral cones
Alternatively, we use the similar strategy in Algorithm 4, the Barzilai-Borwein descent direction with appropriate base is defined as the minimizer of
@@
By optimality conditions, the minimizer of (20) can be written as
@@
The dual problem (DP) is lower dimensional quadratic programming with unit simplex constraint (the vertices of unit simplex constraint are known), then it can be solved by Frank-Wolfe/conditional gradient method efficiently [see, e.g., 41, 7]. However, dual problem of (18) reads as
@@
If $K=\mathbb{R}^{m}_{+}$ and $A=I_{m}$, the subproblem (20) reduces to that of the BBDMO [7]. Consequently, transforming (18) into (20) gives a new insight into BBDMO. Comparing the subproblems (16) and (19), we conclude that the differences between the directions of the steepest descent and the Barzilai-Borwein descent lie in the choice of base of $K^{*}$.
@@ 4       Update $d^{k}$ as the minimizer of (20)
6             return $C$-stationary point $x^{k}$
@@ 9             $x^{k+1}:=x^{k}+t_{k}d^{k}$
Algorithm 8 K-Barzilai-Borwein_descent_method_for_VOPs_with_polyhedral_cones
@@
    $\left\lVert{x^{k+1}-x^{*}}\right\rVert\leq\sqrt{1-\min_{i\in[l]}{t_{k}\left\langle{A_{i},\bm{\mu}}\right\rangle}/{\alpha^{k}_{i}}}\left\lVert{x^{k}-x^{*}}\right\rVert,~{}\forall k\geq 0$.
@@
The following results show that the convergence rate of Algorithm 8 is not sensitive to the choice of transform matrix. More specifically, the descent direction $d^{k}$ and stepsize $t_{k}$ of Algorithm 8 are invariant for some $A\in\mathcal{A}$.
@@
This, together with the reformulated subproblem, implies that $d^{k}_{1}=d^{k}_{2}$. Therefore, $t_{k}^{1}=t_{k}^{2}$ is a consequence of (23). Next, we prove that assertion (23) holds. For any $i\in[l]$, it follows by (22) that there exist $j\in[l]$ such that $A^{1}_{i}=a_{j}A^{2}_{j}$. Notice that $\alpha_{\min}<\alpha_{i}^{k,1},\alpha_{i}^{k,2}<\alpha_{\max}$, $i\in[l]$, without loss of generality, we assume that $\alpha_{i}^{k,1}={\left\lVert{A^{1}_{i}(JF(x^{k})-JF(x^{k-1}))}\right\rVert}/{\left\lVert{s_{k-1}}\right\rVert}$ and $\alpha_{j}^{k,2}={\left\lVert{A^{2}_{j}(JF(x^{k})-JF(x^{k-1}))}\right\rVert}/{\left\lVert{s_{k-1}}\right\rVert}$. Then,
$\frac{A^{1}_{i}}{\alpha_{i}^{k,1}}=\frac{A^{1}_{i}\left\lVert{s_{k-1}}\right\rVert}{\left\lVert{A^{1}_{i}(JF(x^{k})-JF(x^{k-1}))}\right\rVert}=\frac{A^{2}_{j}\left\lVert{s_{k-1}}\right\rVert}{\left\lVert{A^{2}_{j}(JF(x^{k})-JF(x^{k-1}))}\right\rVert}=\frac{A^{2}_{j}}{\alpha_{j}^{k,2}},$
where the second equality is due to the fact that $A^{1}_{i}=a_{j}A^{2}_{j}$. Thus, we have
@@
Assume $\hat{\mathcal{A}}\subset\mathcal{A}$ is bounded and $A_{i}$ is bounded away from 00 for all $A\in\hat{\mathcal{A}}$. Then, there exists $\alpha_{\min}$ and $\alpha_{\max}$ such that the assumption $\alpha_{\min}<\alpha_{i}^{k}<\alpha_{\max}$, $i\in[l]$ holds for all $A\in\hat{\mathcal{A}}$ with $\left\langle{s_{k-1},{A_{i}(JF(x^{k})-JF(x^{k-1}))}}\right\rangle\neq 0$. For the case with linear objective, $\left\langle{s_{k-1},{A_{i}(JF(x^{k})-JF(x^{k-1}))}}\right\rangle=0$ may hold. As illustrated in [7, Example 3], for any $A_{1},A_{2}\in\hat{\mathcal{A}}$, we have $d^{k}_{1}\approx d^{k}_{2}$ with sufficient small $\alpha_{\min}$. As a result, we conclude that the performance of Algorithm 8 is not sensitive to the choice of $A$.
7 Second-order methods from a majorization-minimization perspective
Let $K=\mathbb{R}^{m}_{+}$, Fliege et al. [16] proposed a Newton method for MOPs, the Newton direction is defined as the optimal solution of
$\min\limits_{d\in\mathbb{R}^{n}}\max\limits_{\lambda\in\Delta_{m}}\left\langle{\lambda,JF(x^{k})d+\frac{1}{2}d^{T}D^{2}F(x^{k})d}\right\rangle,$ (24)
where $d^{T}D^{2}F(x^{k})d:=(d^{T}\nabla^{2}F_{1}(x^{k})d,...,d^{T}\nabla^{2}F_{m}(x^{k})d)$. The objective-wise quadratic approximation serves as good surrogate in Newton method but makes the subproblem challenging to solve. From a majorization-minimization perspective, the surrogate should strike a balance between better approximation and lower computational cost. In the remainder of this section, we develop an efficient second-order method. Interestingly, the proposed method coincides with the preconditioned Barzilai-Borwein method for MOPs [9].
Firstly, we note that the challenging nature of the subproblem (24) arises primarily from selecting $\Delta_{m}$ as the base for +$\mathbb{R}^{m}_{+}$. By instead choosing a general base in the subproblem:
$\min\limits_{d\in\mathbb{R}^{n}}\max\limits_{\lambda\in\Delta^{\alpha^{k}}_{m}}\left\langle{\lambda,JF(x^{k})d+\frac{1}{2}d^{T}D^{2}F(x^{k})d}\right\rangle,$ (25)
where $\Delta^{\alpha^{k}}_{m}:=\{c^{*}\in\mathbb{R}^{m}_{+}:\left\langle{c^{*},\alpha^{k}}\right\rangle=1\}$, it can be reformulated as follows:
$\min\limits_{d\in\mathbb{R}^{n}}\max\limits_{\lambda\in\Delta_{m}}\sum\limits_{i\in[m]}\lambda_{i}\frac{\left\langle{\nabla F_{i}(x^{k}),d}\right\rangle}{\alpha^{k}_{i}}+\frac{1}{2}\sum\limits_{i\in[m]}\lambda_{i}\frac{\left\langle{d,\nabla^{2}F_{i}(x^{k})d}\right\rangle}{\alpha^{k}_{i}}.$ (26)
This raises a direct question: how to select an appropriate $\alpha^{k}$? Intuitively, if
$\frac{\left\langle{d^{k},\nabla^{2}F_{i}(x^{k})d^{k}}\right\rangle}{\alpha^{k}_{i}}=\sum\limits_{i\in[m]}\lambda^{k}_{i}\frac{\left\langle{d^{k},\nabla^{2}F_{i}(x^{k})d^{k}}\right\rangle}{\alpha^{k}_{i}}$
for $i\in[m]$, by setting $B_{k}\approx\sum_{i\in[m]}\lambda^{k}_{i}{\nabla^{2}F_{i}(x^{k})}/{\alpha^{k}_{i}}$, we reformulate (26) into the following subproblem:
$\min\limits_{d\in\mathbb{R}^{n}}\max\limits_{\lambda\in\Delta_{m}}\sum\limits_{i\in[m]}\lambda_{i}\frac{\left\langle{\nabla F_{i}(x^{k}),d}\right\rangle}{\alpha^{k}_{i}}+\frac{1}{2}\left\lVert{d}\right\rVert^{2}_{B_{k}},$ (27)
where $\alpha^{k}_{i}\approx\left\langle{d^{k},\nabla^{2}F_{i}(x^{k})d^{k}}\right\rangle/\left\lVert{d^{k}}\right\rVert^{2}_{B_{k}}$ for $i\in[m]$. Since $d^{k}$ is unknown before solving the subproblem, we estimate it with $d^{k-1}$ and set

### v3

@@ 4.2 Improved $K$-steepest descent method for VOPs without line search
As detailed in Remark 4.2, the linear convergence rate can be very slow with imbalanced objectives, this is mainly due to the large gap between $F(\cdot)-F(x^{k})$ and $G_{k,Le}(\cdot)$ from a majorization-minimization perspective. To reduce this gap, one natural strategy is to construct a tighter surrogate function that better approximates the local behavior of $F(\cdot)-F(x^{k})$. Notice that $\bm{\ell}\preceq_{K}L_{\max}e$, we denote the following tighter majorizing surrogate:
@@
[…]
@@
By using the tighter surrogate, we devise the following improved $K$-steepest descent method with coupled subproblems for VOPs.
@@
Algorithm 3 improved $K$-steepest descent method for VOPs with coupled subproblems
@@
If $K=\mathbb{R}^{m}_{+}$, and $e=\mathbf{1}_{m}$, the convergence rate in Lemma 4.2 reduces to that of [8, Corollary 4.3] with $g(\cdot)=0$. Notice that $1/\kappa_{F,\preceq_{K}}\geq\mu_{\min}/L$, which indicates that Algorithm 3 enjoys faster linear convergence than Algorithm 2. Furthermore, by Remark 3.4, we conclude that the improved linear convergence does not depend on the choice of $C_{e}$.
4.3 Trade-off between surrogate gap and per-iteration cost
Although Algorithms 3 exhibits improved linear convergence by using a tighter surrogate function, the per-iteration cost is more expensive than that of Algorithm 2 due to coupled subproblems. Details on solving these subproblems will be provided in Section 6 (see Remark 6.2). In the spirit of majorization-minimization optimization, a direct question arises: how to strike a better trade-off in terms of surrogate gap and per-iteration cost?
Recall that the linear convergence rate of Algorithm 3 does not depend on the choice of $C_{e}$, which is mainly due to coupled subproblems. By denoting
$C_{\bm{\ell}}:=\{c^{*}\in K^{*}:\left\langle{c^{*},\bm{\ell}}\right\rangle=1\},$
we propose the following improved $K$-steepest descent method with seperate subproblems for VOPs.
@@ 4.3 Trade-off between surrogate gap and per-iteration cost
Algorithm 4 improved $K$-steepest descent method for VOPs with seperate subproblems
Using the definition of $C_{\bm{\ell}}$, the seperate subproblem in Algorithm 4 can be rewritten equivalently as the following coupled form:
@@
Therefore, Algorithm 4 enjoys the same improved linear convergence as that of Algorithm 3.
@@
If $K=\mathbb{R}^{m}_{+}$, and $\Delta_{m}^{\bm{\ell}}:=\{c^{*}\in\mathbb{R}^{m}_{++}:\left\langle{c^{*},\bm{\ell}}\right\rangle=1\}$, the Algorithm 4 reduces to [8, Algorithm 5] with $g(\cdot)=0$. Interestingly, the relations between Algorithms 2, 3 and 4 depend solely on the choice of $C_{e}$. If $C_{e}=C_{\bm{\ell}}$, Algorithms 2, 3 and 4 are equivalent. Consequently, regarding the open problem of selecting a better $C_{e}$ mentioned in Remark 4.2, we provide a theoretical answer by setting $C_{e}=C_{\bm{\ell}}$. Furthermore, we can summarize that choosing a tighter surrogate function is equivalent to selecting an appropriate base in the seperate subproblem.
@@
Although Algorithms 3 and 4 both exhibit similar improved linear convergence, the computational cost of solving seperate subproblems is generally lower in Algorithm 4. Details on solving these subproblems will be provided in Section 6 (see Remark 6.2).
5 First-order methods for VOPs with non-majorizing surrogate functions
@@ 5.1 $K$-steepest descent method with line search
Firstly, we revisit $K$-steepest descent method for VOPs with line search.
@@
4    return $K$-stationary point $x^{k}$
@@ 7    $x^{k+1}:=x^{k}+t_{k}d^{k}$
Algorithm 5 K-steepest descent method for VOPs with line search
@@
The following results show that $G_{k,e/t_{\min}}(\cdot)$ is a non-majorizing surrogate function of $F(\cdot)-F(x^{k})$ near $x^{k}$.
@@
By the line search condition, we have
$F(x^{k+1})-F(x^{k})\preceq_{K}JF(x^{k})(x^{k+1}-x^{k})+\frac{1}{2t_{k}}\left\lVert{x^{k+1}-x^{k}}\right\rVert^{2}e\preceq_{K}JF(x^{k})(x^{k+1}-x^{k})+\frac{1}{2t_{\min}}\left\lVert{x^{k+1}-x^{k}}\right\rVert^{2}e.$
Then, the assertion (i) holds. The assertions (ii) and (iii) can be obtained by using the similar arguments as in the proof of Proposition 4.1. ∎
@@
If $K=\mathbb{R}^{m}_{+}$, and $e=\mathbf{1}_{m}$, i.e., $C_{e}=\Delta_{m}$, the convergence rate in Lemma 5.1(ii) reduces to those established in [16, Theorem 4.2] and [49, Theorem 5.6].
5.2 Generic first-order method for VOPs with line search
To reduce the gap between $F(\cdot)-F(x^{k})$ and $G_{k,e/t_{\min}}(\cdot)$, we select $e_{k}\in{\rm int}(K)$ and devise the following generic first-order method:
@@ 2  Select $e_{k}\in{\rm int}(K)$
5    return $K$-stationary point $x^{k}$
@@ 8    $x^{k+1}:=x^{k}+t_{k}d^{k}$
Algorithm 6 Generic first order method for VOPs with line search
It is worth noting that we don’t specify how to select $e_{k}$ in Algorithm 6. This naturally raises the question: what role does $e_{k}$ play in determining the convergence rate? Firstly, we derive the lower bound of stepsize in each iteration.
@@
    $\{x^{k}\}$ converges to an efficient solution $x^{*}$ of (VOP).
@@
As described in Lemma 5.2(iv), by setting $e_{k}=\bm{\mu}$ or $e_{k}=\bm{\ell}$, we can derive the optimal linear convergence rate for Algorithm 6, and the convergence rate reduces to that of Lemma 4.3(ii) with constant $\gamma$. Intuitively, to explore the local curvature information of $F(\cdot)$, we can devise a tighter local surrogate $G_{k,e_{k}/t_{k}}(\cdot)$ with $\bm{\mu}\preceq_{K}e_{k}\preceq_{K}\bm{\ell}$. In this case, the performance of Algorithm 6 can be further improved by using a tighter local surrogate $G_{k,e_{k}/t_{k}}(\cdot)$.
@@
Remark 5.3.
To reduce per-iteration cost, we set $e=e_{k}$, i.e., $C_{e}=C_{e_{k}}$, so that the coupled subproblem in Algorithm 6 can be reformulated into the following seperate form:
$\min_{d\in\mathbb{R}^{n}}\max_{c^{*}\in C_{e_{k}}}\left\langle{c^{*},JF(x^{k})d}\right\rangle+\frac{1}{2}\|d\|^{2}.$
Hence, we conclude that using a variable $C_{e_{k}}$ serves as an appropriate choice of base in SDVO, which provides a theoretical answer to (Q). For $K=\mathbb{R}^{2}_{+}$, Fig. 1 illustrates the choice of the base under the assumption of strong convexity. It suggests that bases should be adaptively selected from the pink region according to (19).
Figure 1: Illustration of the $C_{e_{k}}$ of $K=\mathbb{R}^{2}_{+}$.
6 First-order methods for VOPs with polyhedral cones
@@ 6.1 Steepest descent method for VOPs with polyhedral cones
By using a transform matrix $A\in\mathcal{A}$, the steepest descent direction subproblem for VOPs with polyhedral cones is formulated as follows:
@@
The complete $K$-steepest descent method for VOPs with polyhedral cones is described as follows:
@@ 3   Update $d^{k}$ as the minimizer of (20)
5    return $K$-stationary point $x^{k}$
@@ 7     Compute the stepsize $t_{k}\in(0,1]$ in the following way:
$t_{k}:=\max\left\{\gamma^{j}:j\in\mathbb{N},A(F(x^{k}+\gamma^{j}d^{k})-F(x^{k}))\preceq\gamma^{j}\left(AJF(x^{k})d^{k}+\frac{1}{2}\left\lVert{d^{k}}\right\rVert^{2}\mathbf{1}_{l}\right)\right\}$
$x^{k+1}:=x^{k}+t_{k}d^{k}$
@@
Algorithm 7 K-steepest descent method for VOPs with polyhedral cones
We consider the following surrogate:
$G_{k,A,\mathbf{1}_{l}/t_{k}}(x):=AJF(x^{k})(x-x^{k})+\frac{1}{2t_{k}}\left\lVert{x-x^{k}}\right\rVert^{2}\mathbf{1}_{l}.$ (21)
We can show that $G_{k,A,\mathbf{1}_{l}/t_{k}}(\cdot)$ is a non-majorizing surrogate function of $A(F(\cdot)-F(x^{k}))$ near $x^{k}$.
Proposition 6.1.
Let $G_{k,A,\mathbf{1}_{l}/t_{k}}(\cdot)$ be defined as (21). Then, the following statements hold.
  • (i)
    $A(F(x^{k+1})-F(x^{k}))\preceq G_{k,A,\mathbf{1}_{l}/t_{k}}(x^{k+1})$.
  • (ii)
    If $F(\cdot)$ is $K$-convex, then $G_{k,A,\mathbf{1}_{l}/t_{k}}(\cdot)\in\mathcal{S}_{\mathbf{1}_{l}/t_{k},\mathbf{1}_{l}/t_{k}}(AF,x^{k})$.
  • (iii)
    If $F(\cdot)$ is strongly $K$-convex with $\bm{\mu}\in{\rm int}(K)$, then $G_{k,A,\mathbf{1}_{l}/t_{k}}(\cdot)\in\mathcal{S}_{\mathbf{1}_{l}/t_{k}-A\bm{\mu},\mathbf{1}_{l}/t_{k}}(AF,x^{k})$.
Proof.
@@
    $t_{k}\geq\min\{\min_{i\in[l]}\{\gamma/\left\langle{A_{i},\bm{\ell}}\right\rangle\},1\}$;
@@
    $\left\lVert{x^{k+1}-x^{*}}\right\rVert\leq\sqrt{1-t_{k}\min_{i\in[l]}\left\langle{A_{i},\bm{\mu}}\right\rangle}\left\lVert{x^{k}-x^{*}}\right\rVert,~\forall k\geq 0$.
@@ 6.2 Barzilai-Borwein descent method for VOPs with polyhedral cones
Alternatively, we use the similar strategy in Algorithm 4, the Barzilai-Borwein descent direction subproblem can be rewritten equivalently as the following coupled form:
@@
By Sion’s minimax theorem [43], the minimizer of (25) can be written as
@@
In general, the dual problem (DP) is a lower dimensional quadratic programming with unit simplex constraint (the vertices of unit simplex constraint are known), then it can be solved by Frank-Wolfe/conditional gradient method efficiently [see, e.g., 42, 7]. However, dual problem of (23) reads as
@@
If $K=\mathbb{R}^{m}_{+}$ and $A=I_{m}$, the subproblem (25) reduces to that of the BBDMO [7]. Consequently, transforming (23) into (25) gives a new insight into BBDMO from a majorization-minimization perspective. Comparing the subproblems (20) and (24), it turns out that the differences between the directions of the steepest descent and the Barzilai-Borwein descent lie in the choice of base for $K^{*}$.
@@ 5   Update $d^{k}$ as the minimizer of (25)
7    return $K$-stationary point $x^{k}$
@@ 10    $x^{k+1}:=x^{k}+t_{k}d^{k}$
Algorithm 8 K-Barzilai-Borwein descent method for VOPs with polyhedral cones
We consider the following surrogate:
$G_{k,A,\alpha^{k}/t_{k}}(x):=AJF(x^{k})(x-x^{k})+\frac{1}{2t_{k}}\left\lVert{x-x^{k}}\right\rVert^{2}\alpha^{k}.$ (26)
We can show that $G_{k,A,\alpha^{k}/t_{k}}(\cdot)$ is a non-majorizing surrogate function of $A(F(\cdot)-F(x^{k}))$ near $x^{k}$.
Proposition 6.2.
Let $G_{k,A,\alpha^{k}/t_{k}}(\cdot)$ be defined as (26). Then, the following statements hold.
  • (i)
    $A(F(x^{k+1})-F(x^{k}))\preceq G_{k,A,\alpha^{k}/t_{k}}(x^{k+1})$.
  • (ii)
    If $F(\cdot)$ is $K$-convex, then $G_{k,A,\alpha^{k}/t_{k}}(\cdot)\in\mathcal{S}_{\alpha^{k}/t_{k},\alpha^{k}/t_{k}}(AF,x^{k})$.
  • (iii)
    If $F(\cdot)$ is strongly $K$-convex with $\bm{\mu}\in{\rm int}(K)$, then $G_{k,A,\alpha^{k}/t_{k}}(\cdot)\in\mathcal{S}_{\alpha^{k}/t_{k}-A\bm{\mu},\alpha^{k}/t_{k}}(AF,x^{k})$.
Proof.
@@
    $\left\lVert{x^{k+1}-x^{*}}\right\rVert\leq\sqrt{1-t_{k}\min_{i\in[l]}\{{\left\langle{A_{i},\bm{\mu}}\right\rangle}/{\alpha^{k}_{i}}\}}\left\lVert{x^{k}-x^{*}}\right\rVert,~\forall k\geq 0$.
@@
The following result shows that the convergence rate of Algorithm 8 is not sensitive to the choice of transform matrix. More specifically, the descent direction $d^{k}$ and stepsize $t_{k}$ of Algorithm 8 are invariant for some $A\in\mathcal{A}$.
@@
This, together with the reformulated subproblem, implies that $d^{k}_{1}=d^{k}_{2}$. Therefore, $t_{k}^{1}=t_{k}^{2}$ is a consequence of (29).
Next, we prove that assertion (29) holds. For any $i\in[l]$, it follows by (28) that there exist $j\in[l]$ such that $A^{1}_{i}=a_{j}A^{2}_{j}$. Notice that $\alpha_{\min}<\alpha_{i}^{k,1},\alpha_{i}^{k,2}<\alpha_{\max}$, $i\in[l]$, we distinguish two cases:
$\alpha_{i}^{k,1}={\left\lVert{A^{1}_{i}(JF(x^{k})-JF(x^{k-1}))}\right\rVert}/{\left\lVert{s_{k-1}}\right\rVert}$ and $\alpha_{j}^{k,2}={\left\lVert{A^{2}_{j}(JF(x^{k})-JF(x^{k-1}))}\right\rVert}/{\left\lVert{s_{k-1}}\right\rVert}$,
and
$\alpha_{i}^{k,1}={\left\langle{A^{1}_{i}(JF(x^{k})-JF(x^{k-1})),s_{k-1}}\right\rangle}/{\left\lVert{s_{k-1}}\right\rVert^{2}}$ and $\alpha_{i}^{k,2}={\left\langle{A^{1}_{i}(JF(x^{k})-JF(x^{k-1})),s_{k-1}}\right\rangle}/{\left\lVert{s_{k-1}}\right\rVert^{2}}$.
Since $A^{1}_{i}=a_{j}A^{2}_{j}$ ($a_{j}>0$), it is easy to verify that
$\frac{A^{1}_{i}}{\alpha_{i}^{k,1}}=\frac{A^{2}_{j}}{\alpha_{j}^{k,2}}$
holds in both cases. Thus, we have
@@
Assume $\hat{\mathcal{A}}\subset\mathcal{A}$ is bounded and $A_{i}$ is bounded away from 0 for all $A\in\hat{\mathcal{A}}$. Then, there exists $\alpha_{\min}$ and $\alpha_{\max}$ such that the assumption $\alpha_{\min}<\alpha_{i}^{k}<\alpha_{\max}$, $i\in[l]$ holds for all $A\in\hat{\mathcal{A}}$ with $\left\langle{s_{k-1},{A_{i}(JF(x^{k})-JF(x^{k-1}))}}\right\rangle\neq 0$. For the case with linear objective, $\left\langle{s_{k-1},{A_{i}(JF(x^{k})-JF(x^{k-1}))}}\right\rangle=0$ may hold. As illustrated in [7, Example 3], for any $A_{1},A_{2}\in\hat{\mathcal{A}}$, we have $d^{k}_{1}\approx d^{k}_{2}$ with sufficient small $\alpha_{\min}$. As a result, we conclude that the performance of Algorithm 8 is not sensitive to the choice of $A$.
6.3 Backtracking method for VOPs with polyhedral cones
As described in Lemma 6.2, we apply Barzilai-Borwein method to ensure $A\bm{\mu}\preceq\alpha^{k}\preceq A\bm{\ell}$, which in turn is expected to obtain $A\bm{\mu}\preceq\alpha^{k}/t_{k}\preceq A\bm{\ell}$, thereby achieving a fast linear convergence rate in practice. However, this strategy actually yields an even slower convergence rate.
Proposition 6.4.
Assume that $F(\cdot)$ is strongly $K$-convex with $\bm{\mu}\in{\rm int}(K)$, where $K=\{x\in\mathbb{R}^{m}:0\preceq Ax\}.$ Let $\{x^{k}\}$ be the sequence generated by Algorithm 8 and $x^{*}$ be the efficient solution satisfies $F(x^{*})\preceq_{K}F(x^{k})$ for all $k\geq 0$. Then